OpenStudy (anonymous):

substitution: the integral of cos3xdx, u=3x

OpenStudy (anonymous):

cosu du/3

myininaya (myininaya):

\[du=3 d\[\frac{1}{3}\int\limits_{}^{}\cos(u) du=\frac{1}{3}\sin(u)+C\]x\]

OpenStudy (anonymous):

-sinu/3+c

myininaya (myininaya):

not negative derivative of sinu is cosu

OpenStudy (anonymous):

u = 3x, du = 3dx \[\int(cos(3x)dx=\frac{1}{3}\int cos(u)u =\frac{1}{3}sin(u)=\frac{1}{3}sin(3x)\]

OpenStudy (anonymous):

ohh sorry

myininaya (myininaya):

=1/3*sin(3x)+C

OpenStudy (anonymous):

oh yeah! +C

myininaya (myininaya):

satellite your answer was so far off without that C

myininaya (myininaya):

lol

OpenStudy (anonymous):

ok so who's right?

OpenStudy (anonymous):

obviously me

OpenStudy (anonymous):

remember, i'm the one that doesn't know what's going on :]

OpenStudy (anonymous):

oh, right obviouslyyy. silly me.